Forrester Fashions has annual credit sales of 250,000 units with an average collection period of 70 days. The company has a per-unit variable cost of $20 and a per-unit sale price of $30. Bad debts currently are 5% of sales. The firm estimates that a proposed relaxation of credit standards would not affect its 70-day average collection period but would increase bad debts to 7.5% of sales, which would increase to 300,000 units per year. Forrester requires a 12% return on investments. Show all necessary calculations required to evaluate Forrester
Answer #1
Ans. Ans. Existing New
Sales revenue (250000 x 30) = 7500000 9000000
Variable cost (250000 x 20) = -5000000 6000000
Contribution 2500000 3000000
Bad debts -375000 -675000
Interest -175000 -210000
Total 1950000 2115000
As the new proposed policy has increased the profit of the firm it should beaccepted.
Answer #2

Increase in Account Receivable
50000*30/365*70= $287671

Increase in bad debts Bad debts New Sales 9000000 675000 Old Sales 7500000 375000 300000

Increase in Contribution Margin = 50000*30-20 = 500000

Less increase in bad debts = -300000

Less opportunity cost of Account Rcceivable = 287671*.12 = -34521

Net benefit of relaxation =165479

Yes the policy should be relaxed
Answer #3

SEE THE IMAGE. ANY DOUBTS, FEEL FREE TO ASK. THUMBS UP PLEASE